OpenStudy (anonymous):

Arianna kicks a soccer ball off the ground and into the air with an initial velocity of 42 feet per second. Assume the starting height of the ball is 0 feet. Approximately, how long does it take until the soccer ball hits the ground again?

OpenStudy (anonymous):

1.3 secs 2.6 secs 16.0 secs 27.6 secs

OpenStudy (anonymous):

What is the angle?

OpenStudy (anonymous):

its math or physics

OpenStudy (anonymous):

physics

OpenStudy (anonymous):

@allie_bear22 what is the angle???

OpenStudy (anonymous):

im not sure

OpenStudy (anonymous):

and its math

OpenStudy (anonymous):

@lordcyborg

OpenStudy (anonymous):

ya wait i'm trying to figure out the angle

OpenStudy (anonymous):

i think that the angle is vertical

OpenStudy (anonymous):

if u take it to be vertical throw then answer comes out to be 8.4 sec

OpenStudy (anonymous):

@allie_bear2 This question is from which topic???

OpenStudy (anonymous):

but thats not an option so maybe its a diffrent angle

OpenStudy (anonymous):

ya wt is the related topic???

OpenStudy (anonymous):

math

OpenStudy (anonymous):

i think you must use this formula h = -16t^2 + Vot + s

OpenStudy (anonymous):

so initial velocity of 42 feet per second = Vo=42 Assume the starting height of the ball is 0 feet = s=0

OpenStudy (anonymous):

@allie_bear22

OpenStudy (anonymous):

so its 26?

OpenStudy (anonymous):

h = -16t^2 + 42t the soccer ball hits the ground again => h=0 so now put zero for h 0=-16t^2+42t

OpenStudy (anonymous):

so next step -16t^2 + 42t = 0 t(-16t+42) = 0 t = 42/16 = 2.625 sec

OpenStudy (anonymous):

so the answer is 2.6

OpenStudy (anonymous):

thnk yu

OpenStudy (anonymous):

YW!